# Axi-majoron for almost everything

### v1

@@ Abstract
The details of the minimal cosmological standard model (MCSM) proposed in Ref. [1] are discussed. The model is based on the scale-symmetry and the global Peccei-Quinn(PQ) symmetry with a key assumption that the latter is broken only in the gravity sector in a scale-invariant manner. We show that the model provides a quite simple unified framework for the unknown history of the universe from inflation to the epoch of big-bang nucleosynthesis, simultaneously addressing key puzzles of high energy theory and cosmology: (i) the origin of scales, (ii) primordial inflation, (iii) matter-antimatter asymmetry, (iv) tiny neutrino masses, (v) dark matter, and (vi) the strong CP-problem. Scale symmetry can be exact, and the Planck scale is dynamically generated. The presence of Gauss-Bonnet term may safely retain dangerous non-perturbative symmetry-breaking effects negligible, allowing large-field trans-Planckian inflation along the PQ-field. Iso-curvature perturbations of axi-majorons are suppressed. A sizable amount of PQ-number asymmetry is generated at the end of inflation, and conserved afterwards. Domain wall problem is absent due to the PQ-number asymmetry. Baryogenesis can be realized by either inflationary Affleck-Dine mechanism or spontaneous leptogenesis thanks to the PQ-number asymmetry, or by resonant leptogenesis. Dark matter can be purely cold axi-majorons from the mis-alignment contribution only with the symmetry-breaking scale of $\mathcal{O}(10^{12}){\rm GeV}$. Hot axi-majorons from the decay of the inflaton become a natural source for a sizable amount of dark radiation. Inflationary gravitational waves are expected to have information about some masse parameters of the left-handed and the right-handed neutrinos, thanks to the presence of an early matter-domination era driven by at least a long-lived right-handed neutrino species.
@@ 1 Introduction
Recently, there have been several works discussing dynamical generation of scales in a scale invariant version of gravity and extensions of the SM [5, 6, 7, 8]. An interesting finding is that the Planck scale can be dynamically generated even without a Coleman-Weinberg(CW)-type symmetry-breaking [9], thanks to the conservation of scale current which has a kernel settling down to a constant [5, 8]. Low energy scales, such as the electroweak scale, can be generated in connection with the field responsible for the generation of the Planck scale, or separately via CW-mechanism in the presence of matter fields other than SM fields. Scale-invariance can be maintained even at quantum level, provided that the renormalization scale (which usually appears as an explicit dimensionful parameter) is replaced by a fundamental scalar field. Preservation of scale-invariance decouples the massless dilaton from the other fields, hence makes the theory free from fifth-force constraints [10, 11]. This implies that a fully scale-invariant theory can be consistent with low energy phenomenology.
@@
Meanwhile, any theory trying to explain the current state of the universe should be able to accommodate a phase of primordial inflation  [^2] It was found that in the scenario of an emergent gravity with scale-invariance a slow-roll inflation can arise quite naturally [7, 12]. However, connection with low energy physics was missing., generation of matter-antimatter asymmetry, proper candidate(s) of dark matter, neutrino masses inferred by neutrino flavor oscillation data [13], and probably a solution to the strong CP-problem such as invisible axion [14, 15, 16, 17]. All of these big puzzles of cosmology and phenomenology can not be solved within the SM only, and definitely demand a theory beyond the standard model(BSM). In this regard, the idea of unifying axion and majoron [18] is quite interesting, since it can be realized in a scale-invariant manner, simultaneously address the tiny neutrino masses, and provide an axion solution to the strong CP-problem in a very simple structure. Then, it is tempting to see if the idea can be combined with inflation and generation of matter-antimatter asymmetry while keeping the scale-invariance.
@@
In Ref. [1], we proposed a minimal scale-invariant $U(1)_{\rm PQ}$-symmetric extension of the SM [19, 20], dubbed as ‘the minimal cosmological standard model(MCSM)’, to address the aforementioned issues altogether other than the notorious Cosmological Constant(CC) problem. In this work, we show the details of the MCSM. It is a DFSZ-type axion model [16, 17] unified with the singlet majoron model, in which Peccei-Quinn(PQ) field not only provides the axion solution, but it is also responsible for the generation of the masses of right-handed neutrinos. Hence, we call the associated pseudo-Goldstone boson as ‘axi-majoron’. In the sense of unifying axion and majoron, our scenario is similar to the earlier works such as Refs. [18] and [21, 22], based on KSVZ [14, 15] and DFSZ [16, 17] axion models, respectively. Also, in regard of providing a unified framework addressing aforementioned puzzles, one can find works such as Refs. [23, 24]. Our scenario is different from those works in various aspects including dynamical generation of the Planck scale, inflationary physics, generation of matter-antimatter asymmetry and physics of axions. In particular, assuming $U(1)_{\rm PQ}$-symmetry to be broken in the gravity sector only [^3] There are some earlier works such as Refs. [25] and [26] where the PQ symmetry-breaking non-minimal gravitational interaction was considered. In Ref. [25], the authors payed attention only to the impact of such an interaction on the axion-isocurvature perturbations which can be suppressed by a temporary axion mass much larger than the expansion rate during inflation thanks to the interaction. In Ref. [26], angular motion was removed by setting initial position of the field at the valley of the angular potential for simplicity. In both works, the PQ-field are located at the angular minimum position during inflation. In our case, the angular motion and the non-zero angular displacement of the PQ-field at the end of inflation play the critical role for the post inflation cosmology., we show that the dynamics of the inflaton along axi-majoron direction can generate a large amount of PQ-number asymmetry which is eventually converted to the baryon number asymmetry via the seesaw sector. It is a kind of Affleck-Dine(AD) mechanism [27] realized by Peccei-Quinn field. Interestingly, such an asymmetry corresponding to a motion of axi-majoron allows not only AD-mechanism but also spontaneous leptogenesis. Even in the case of PQ-number asymmetry not being enough, our scenario allows resonant leptogenesis [28] to work. Also, very intriguingly, in regard to the axion solution to the strong CP-problem, the symmetry-breaking terms in the gravity sector are harmless at least at perturbative level, and may be so even at non-perturbative level thanks to topological curvature-terms such as Gauss-Bonnet [29]. Iso-curvature perturbations are suppressed, and the domain-wall problem is absent due to the angular motion of PQ-field. Also, a sizable amount of dark radiation from hot axi-majorons can be obtained naturally, and may ameliorate the Hubble tension [30] [31]. The standard thermal bath is recovered by decays of at least a long-lived non-relativistic right-handed neutrino species. Because of the presence of an early matter domination era, primordial gravitational waves from inflation are expected to have a characteristic spectral feature, containing information of the masses of the lightest left-handed and the lightest right-handed neutrinos. In combinations with dark radiation and inflationary observables, it may provide a unique way to probe such physical quantities. Our scenario provides thus a very minimal, unique unified framework for the unknown history of the universe from inflation to the BBN epoch.
@@ 2 The Model
Instead of considering all these fields explicitly, we deliver our argument with a simplified version, having only one SM Higgs doublet $H$. The presence of the second Higgs doublet which is essential for a realization of DFSZ axion scenario does not affect the main argument. Paying attention to only these fields and their relevant interactions, we characterize our model by the following action555 The largrangian of the gravity sector can have terms involving $H_{1}H_{2}$ such as $\tilde{R}\left(H_{1}^{\dagger}H_{2}\pm{\rm h.c.}\right)$. Also, in the matter sector, there can be terms such as $\Phi^{2}\left(H_{1}^{\dagger}H_{2}\right)+{\rm h.c.}$. We ignore these terms in this work since we will be interested in the case in which the couplings of $H_{1}$ and $H_{2}$ to $\Phi$ are negligible and hence those Higgs fields are likely to stay at the symmetric phase before the electroweak phase transition (EWPT). We will discuss else where the case where initially $H_{1}$ and $H_{2}$ have very large field values during inflation. :
@@ 3.3 Reheating after inflation
From the seesaw formula for the mass parameters of the left-handed neutrinos [58, 59]  [^17] At low energy below the electroweak scale, $y_{\nu}$ should be replaced to the one obtained from a bi-unitarity transformation associated with physical mass-eigenstates. We ignore such details here, since it does not make changes of the main point of our argument. Hence, we assumed $y_{\nu}$ is flavor-diagonal for simplicity.,
@@
$V(\phi)\approx\lambda_{\phi}\left[\phi_{0}^{2}{\tilde{\phi}}^{2}+\phi_{0}{\tilde{\phi}}^{3}+\frac{1}{4}{\tilde{\phi}}^{4}\right]\ \xrightarrow[]{{\tilde{\phi}}\ll\phi_{0}}\ \frac{1}{2}m_{\phi}^{2}{\tilde{\phi}}^{2}\ .$ (3.64)
The energy density of $\phi$ is then given by
$\rho_{\phi}\approx\frac{1}{2}m_{\phi}^{2}{\tilde{\phi}}^{2}\ ,$ (3.65)
where ${\tilde{\phi}}$ is understood as the amplitude of oscillation, and $m_{\phi}=\sqrt{2\lambda_{\phi}}\phi_{0}$. In this regime, the universe is in a matter-domination era driven by non-relativistic massive $\phi$-particles [^18] There is no kination era driven by the kinetic anergy caused by angular motion of $\Phi$. This is because the phase velocity of $\Phi$, i.e., $\dot{\theta}$ scales as $a^{-1}$ for $\phi\gtrsim\phi_{\times}$ and $a^{-3}$ for $\phi\lesssim\phi_{\times}$ in our scenario. Its energy contribution is always sub-dominant relative to the contribution from the radial mode.. Hence, requiring $H=(2/3)\Gamma^{\phi}$, one finds the oscillation amplitude ${\tilde{\phi}}_{*}$ at the epoch of the decay of $\phi$ as
@@
$H_{*}\approx\frac{\sqrt{2}\lambda_{\phi}^{3/2}}{24\pi}\phi_{0}\simeq 3.5\times 10^{-4}{\rm GeV}\times\xi_{\phi}^{3}\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ .$ (3.67)
@@
One may worry about the possibility of holding $\phi$ around the origin due to thermal effect before $\phi$ falls below $\phi_{\times}$. If it happens, we may have to worry about domain-wall problem although the details may also depend on angular motion of $\Phi$. Thermal effect would be from the interaction of $\phi$ to RHNs which had been produced already from the partial decay of $\phi$. Those RHNs have energy of order of $m_{\phi}\approx\sqrt{3\lambda_{\phi}}\phi_{\times}$ at the epoch when $\phi\sim\phi_{\times}$ before either thermal or kinetic equilibrium is achieved. The number density is $n_{N_{i}}\approx 2{\rm B}_{i}\left(\Gamma^{\phi}/H_{\times}\right)\rho_{\phi}/m_{\phi}$, and their scattering cross-section to $\phi$ and axi-majorons is $\sigma=c_{\sigma}\times(\lambda_{\phi}^{4}/m_{\phi}^{2})$ with $c_{\sigma}=\mathcal{O}(0.1)$ being a numerical constant. Then, the ratio of the scattering rate $\Gamma_{s}$ to the expansion at the epoch, i.e., $H\sim H_{\times}\equiv H(\phi_{\times})$ with $H$ given in Eq. (3.62), is expected to be
$\frac{\Gamma_{s}}{H_{\times}}\approx\frac{3c_{\sigma}{\rm B}_{i}\lambda_{\phi}^{4}}{32\pi}\left(\frac{M_{\rm P}}{\phi_{0}}\right)^{2}.$ (3.68)
Hence, thermal effect can be ignored as long as the non-minimal gravitational coupling $\xi_{\phi}$ is constrained as
$\xi_{\phi}<\frac{350}{{\rm B}_{i}^{1/8}}\times\left(\frac{\phi_{0}}{10^{9}{\rm GeV}}\right)^{1/4}\ .$ (3.69)
For simplicity, we restrict ourselves to this case in the following discussion.
@@
It should be noted that, even though thermal effect is negligible, domain-wall problem can be avoided only if $\Phi$ has angular motion which is large enough to overcome its quantum fluctuations. Otherwise, $\Phi$ would spread over the whole $2\pi$ angle, causing the domain-wall problem even though it may not get through a phase transition from the origin.
Now, we pay attention to the case of $\phi_{*}<\phi_{\times}$ first, which corresponds to
$\xi_{\phi}<100\times\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)^{1/2}\ ,$ (3.70)
@@
and will comment on the opposite case (i.e., $\phi_{*}\gtrsim\phi_{\times}$) later. Once $\phi$ decays at the epoch of $H_{*}$, the universe will be filled with axi-majorons and RHNs as a comparable or sub-dominant component. Depending on interaction rates, they may or may not reach thermal or kinetic equilibrium among themselves before the standard thermal bath is established from decays of RHNs. Initially, the production of $\phi$ via inverse decay is kinematically forbidden. Hence, thermalization before the decay of RHNs would be driven by interactions of axi-majoron kinetic term, $y_{N_{i}}$, and $y_{\nu_{i}}$. The $\phi$-mediating self-interaction cross-section of axi-majorons is given by
$\sigma_{a,s}=\frac{3}{32\pi s}\left(\frac{m_{\phi}}{\phi_{0}}\right)^{4}\ ,$ (3.71)
where $s=E_{\rm CM}^{2}$ with $E_{\rm CM}$ the total energy at the center of momentum(CM) frame. Then the scattering rate is given by
$\Gamma_{a,s}=\frac{\lambda_{\phi}^{3}E_{a}}{128\pi^{3}}\left(\frac{M_{\rm P}}{\phi_{0}}\right)^{2}\ ,$ (3.72)
where $E_{a}$ is the energy of axi-majoron in the CM frame. The expansion rate of the universe during radiation-domination era scales as
$H(E_{a})=H_{*}\left(\frac{2E_{a}}{m_{\phi}}\right)^{2}\ .$ (3.73)
From the condition, $H\lesssim\Gamma_{a,s}/2$ for equilibration, the expansion rate at the epoch of axi-majoron kinetic-equilibrium among themselves is then given by
$H_{a,\rm equil}$ $=$ $H_{*}\times\left[\frac{\sqrt{3}\lambda_{\phi}}{8\pi}\left(\frac{M_{\rm P}}{\phi_{0}}\right)\right]^{4}$ (3.74)
$\simeq$ $H_{*}\times\left(\frac{\xi_{\phi}}{10^{2}}\right)^{8}\left(\frac{\phi_{0}^{\rm ref}}{\phi_{0}}\right)^{4}\ .$
@@
where $N=\left(\nu_{R},\nu_{R}^{c}\right)^{T}$ is a four-component Majorana field with the flavor index suppressed, and only relevant leading order terms are shown. When the condition in Eq. (3.56) is satisfied for the RHN-channels of the decay of the inflaton, the cross-sections of the various interaction between the axi-majorons and the RHNs are smaller than the one for axi-majoron self-interactions by a factor of $\mathcal{O}(0.1-1)\times b_{i}^{2}$. On the other hand, from the ratio of $y_{\nu_{i}}$ to $y_{N_{i}}$, which is given by
@@
one may think that the interaction $y_{\nu_{i}}$ can cause much more efficient scatterings and annihilations of $N_{i}$s to SM particles. However, the decay rate of a RHN mass eigenstate $N_{i}$ is given by (see Appendix C)
@@
where $m_{N_{i}}\gg m_{h}$ was assumed. Also, if the left-handed neutrino mass-eigenstates have the normal hierarchy [13] for their masses as an example, $m_{\nu_{i}}\approx m_{\nu}$ is required for heavy RHN mass-eigenstates $N_{2,3}$. Then, for $\xi_{\phi}$ in the range we are interested in, with $\phi_{0}\sim\phi_{0}^{\rm ref}$, heavy RHN mass-eigenstates decay well before their self-interactions mediated by either axi-majorons or SM particles become relevant. The lightest RHNs can also be decoupled from axi-majorons and SM particles before their decay if $\xi_{\phi}\ll\mathcal{O}(10^{2})$ even though $m_{\nu_{1}}\ll m_{\nu}$ (see Eqs. (3.74) and (3.76)).
Once produced, depending on parameters, axi-majorons and RHNs may or may not evolve independently until the decay of the latter. If a RHN mass-eigenstate evolves independently at least until its decay, it becomes non-relativistic when the expansion rate becomes
$H_{i,\rm non}=H_{*}\left(\frac{2m_{N_{i}}}{m_{\phi}}\right)^{2}=2b_{i}H_{*}\ .$ (3.78)
@@
$\frac{\Gamma_{N_{i}}}{H_{i,\rm non}}=\frac{3}{\sqrt{2\lambda_{\phi}}}\frac{m_{\nu_{i}}\phi_{0}}{v_{h}^{2}}\simeq\frac{66}{\xi_{\phi}}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ .$ (3.79)
Also, if stable enough, it would start dominating the universe when the momentum of axi-majoron becomes $p_{a}=p_{a,\rm eq}^{i}$ with
$p_{a,\rm eq}^{i}$ $=$ $2{\rm B}_{i}m_{i}\ .$ (3.80)

### v2

@@ Abstract
The details of the minimal cosmological standard model (MCSM) proposed in Ref. [1] are discussed. The model is based on the scale-symmetry and the global Peccei-Quinn(PQ) symmetry with a key assumption that the latter is broken only in the gravity sector in a scale-invariant manner. We show that the model provides a quite simple unified framework for the unknown history of the universe from inflation to the epoch of big-bang nucleosynthesis, simultaneously addressing key puzzles of high energy theory and cosmology: (i) the origin of scales, (ii) primordial inflation, (iii) matter-antimatter asymmetry, (iv) tiny neutrino masses, (v) dark matter, and (vi) the strong CP-problem. Scale symmetry can be exact, and the Planck scale is dynamically generated. The presence of Gauss-Bonnet term may safely retain dangerous non-perturbative symmetry-breaking effects negligible, allowing a large-field trans-Planckian inflation along the PQ-field. Iso-curvature perturbations of axi-majorons are suppressed. A sizable amount of PQ-number asymmetry is generated at the end of inflation, and conserved afterwards. Domain wall problem is absent due to the non-restoration of the symmetry and the non-zero PQ-number asymmetry. Baryogenesis can be realized by either the transfer of the PQ-number asymmetry through the seesaw sector, or by resonant leptogenesis. Dark matter is purely cold axi-majorons from the mis-alignment contribution with the symmetry-breaking scale of $\mathcal{O}(10^{12}){\rm GeV}$. Hot axi-majorons from the decay of the inflaton become a natural source for a sizable amount of dark radiation. Inflationary gravitational waves have information about the masse parameters of the lightest left-handed and right-handed neutrinos, thanks to the presence of an early matter-domination era driven by the long-lived lightest right-handed neutrino species.
@@ 1 Introduction
Recently, there have been several works discussing dynamical generation of scales in a scale invariant version of gravity and extensions of the SM [5, 6, 7, 8, 9]. An interesting finding is that the Planck scale can be dynamically generated even without a Coleman-Weinberg(CW)-type symmetry-breaking [10], thanks to the conservation of scale current which has a kernel settling down to a constant [5, 9]. Low energy scales, such as the electroweak scale, can be generated in connection with the field responsible for the generation of the Planck scale, or separately via CW-mechanism in the presence of matter fields other than SM fields. Scale-invariance can be maintained even at quantum level, provided that the renormalization scale (which usually appears as an explicit dimensionful parameter) is replaced by a fundamental scalar field. Preservation of scale-invariance decouples the massless dilaton from the other fields, hence makes the theory free from fifth-force constraints [11, 12]. This implies that a fully scale-invariant theory can be consistent with low energy phenomenology.
@@
Meanwhile, any theory trying to explain the current state of the universe should be able to accommodate a phase of primordial inflation  [^2] It was found that in the scenario of an emergent gravity with scale-invariance a slow-roll inflation can arise quite naturally [5, 13, 8, 14]., generation of matter-antimatter asymmetry, proper candidate(s) of dark matter, neutrino masses inferred by neutrino flavor oscillation data [15], and probably a solution to the strong CP-problem such as invisible axion [16, 17, 18, 19]. All of these big puzzles of cosmology and phenomenology can not be solved within the SM only, and definitely demand a theory beyond the standard model(BSM). In this regard, the idea of unifying axion and majoron [20] is quite interesting, since it can be realized in a scale-invariant manner, simultaneously address the tiny neutrino masses, and provide an axion solution to the strong CP-problem in a very simple structure. Then, it is tempting to see if the idea can be combined with inflation and generation of matter-antimatter asymmetry while keeping the scale-invariance.
@@
In Ref. [1], we proposed a minimal scale-invariant $U(1)_{\rm PQ}$-symmetric extension of the SM [21, 22], dubbed as ‘the minimal cosmological standard model(MCSM)’, to address the aforementioned issues altogether other than the notorious Cosmological Constant(CC) problem. In this work, we show the details of the MCSM. It is a DFSZ-type axion model [18, 19] unified with the singlet majoron model, in which Peccei-Quinn(PQ) field not only provides the axion solution, but it is also responsible for the generation of the masses of right-handed neutrinos. Hence, we call the associated pseudo-Goldstone boson as ‘axi-majoron’. In the sense of unifying axion and majoron, our scenario is similar to the earlier works such as Refs. [20] and [23, 24], based on KSVZ [16, 17] and DFSZ [18, 19] axion models, respectively. Also, in regard of providing a unified framework addressing aforementioned puzzles, one can find works such as Refs. [25, 26]. Our scenario is different from those works in various aspects including dynamical generation of the Planck scale, inflationary physics, generation of matter-antimatter asymmetry and physics of axions. In particular, assuming $U(1)_{\rm PQ}$-symmetry to be broken in the gravity sector only [^3] There are some earlier works such as Refs. [27] and [28] where the PQ symmetry-breaking non-minimal gravitational interaction was considered. In Ref. [27], the authors payed attention only to the impact of such an interaction on the axion-isocurvature perturbations which can be suppressed by a temporary axion mass much larger than the expansion rate during inflation thanks to the interaction. In Ref. [28], initial position of the field was set at the valley of the angular potential for simplicity, removing the impact of angular motion. In both works, the PQ-field are located at the angular minimum position during inflation. In our case, the angular motion and the non-zero angular displacement of the PQ-field at the end of inflation play the critical role for the post inflation cosmology., we show that the dynamics of the inflaton along axi-majoron direction can generate a large amount of PQ-number asymmetry which is eventually converted to the baryon number asymmetry through the seesaw sector. It is a kind of Affleck-Dine(AD) mechanism [29] realized by Peccei-Quinn field. Interestingly, such an asymmetry corresponding to a motion of axi-majoron allows not only AD-mechanism but also spontaneous leptogenesis although the latter turns out to be sub-dominant. Even in the case of PQ-number asymmetry not being enough, our scenario allows resonant leptogenesis [30] to work. Also, very intriguingly, in regard to the axion solution to the strong CP-problem, the symmetry-breaking terms in the gravity sector are harmless at least at perturbative level, and may be so even at non-perturbative level thanks to topological curvature-terms such as Gauss-Bonnet [31]. Iso-curvature perturbations are suppressed, and the domain-wall problem is absent due to the angular motion of PQ-field with non-restoration of the symmetry. Also, a sizable amount of dark radiation from hot axi-majorons can be obtained naturally, and may ameliorate the Hubble tension [32] [33]. The standard thermal bath is recovered by decays of the long-lived lightest right-handed neutrino species. Because of the presence of an early matter domination era, primordial gravitational waves from inflation are expected to have a characteristic spectral feature, containing information of the masses of the lightest left-handed and the lightest right-handed neutrinos. In combinations with dark radiation and inflationary observables, it may provide a unique way to probe such physical quantities. Our scenario provides thus a very minimal, unique unified framework for the unknown history of the universe from inflation to the BBN epoch.
@@ 2 The Model
Instead of considering all these fields explicitly, we deliver our argument with a simplified version, having only one SM Higgs doublet $H$. The presence of the second Higgs doublet which is essential for a realization of DFSZ axion scenario does not affect the main argument. Paying attention to only these fields and their relevant interactions, we characterize our model by the following action555 The largrangian of the gravity sector can have additional symmetry-breaking terms involving $H_{1}^{\dagger}H_{2}$ such as $\tilde{R}\left(H_{1}^{\dagger}H_{2}\pm{\rm h.c.}\right)$. Also, in the matter sector, there can be terms such as $\Phi^{2}\left(H_{1}^{\dagger}H_{2}\right)+{\rm h.c.}$. We ignore these terms in this work since we will be interested in the case in which the couplings of $H_{1}$ and $H_{2}$ to $\Phi$ are negligible and hence those Higgs fields are likely to stay at the symmetric phase before the electroweak phase transition (EWPT). We will discuss else where the case where initially $H_{1}$ and $H_{2}$ have very large field values during inflation. :
@@ 3.3 Reheating after inflation
From the seesaw formula for the mass parameters of the left-handed neutrinos [60, 61]  [^17] At low energy below the electroweak scale, $y_{\nu}$ should be replaced to the one obtained from a bi-unitarity transformation associated with physical mass-eigenstates. We ignore such details here, since it does not make changes of the main point of our argument. In this reason, we assumed $y_{\nu}$ is flavor-diagonal for simplicity.,
@@
$V(\phi)\approx\lambda_{\phi}\left[\phi_{0}^{2}{\tilde{\phi}}^{2}+\phi_{0}{\tilde{\phi}}^{3}+\frac{1}{4}{\tilde{\phi}}^{4}\right]\ \xrightarrow[]{{\tilde{\phi}}\ll\phi_{0}}\ \frac{1}{2}m_{\phi,0}^{2}{\tilde{\phi}}^{2}$ (3.64)
where $m_{\phi,0}\equiv\sqrt{2\lambda_{\phi}}\phi_{0}$. The energy density of $\phi$ is then given by
$\rho_{\phi}\approx\frac{1}{2}m_{\phi,0}^{2}{\tilde{\phi}}^{2}\ ,$ (3.65)
where ${\tilde{\phi}}$ is understood as the amplitude of oscillation. In this regime, the universe is in a matter-domination era driven by non-relativistic massive $\phi$-particles [^18] There is no kination era driven by the kinetic anergy caused by angular motion of $\Phi$. This is because the phase velocity of $\Phi$, i.e., $\dot{\theta}$ scales as $a^{-1}$ for $\phi\gtrsim\phi_{\times}$ and $a^{-3}$ for $\phi\lesssim\phi_{\times}$ in our scenario. Its energy contribution is always sub-dominant relative to the contribution from the radial mode.. Hence, requiring $H=(2/3)\Gamma^{\phi}$, one finds the oscillation amplitude ${\tilde{\phi}}_{*}$ at the epoch of the decay of $\phi$ as
@@
${\tilde{H}}_{*}\approx\frac{\sqrt{2}\lambda_{\phi}^{3/2}}{24\pi}\phi_{0}\simeq 3.5\times 10^{-4}{\rm GeV}\times\xi_{\phi}^{3}\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ .$ (3.67)
@@
One may worry about the possibility of holding $\phi$ around the origin due to thermal effect before $\phi$ falls below $\phi_{\times}$. If it happens, the domain-wall problem can arise although the details may also depend on angular motion of $\Phi$. In the vicinity of the origin, thermal effect is from the interaction of $\Phi$ to RHNs which had been produced already from the partial decay of $\phi$. If those RHNs were thermalized with axi-majorons soon after their production, for $\phi_{*}>\phi_{\times}$, the temperature around the epoch of $\phi$-particles’ decay is expected to be
$T_{N,*}\approx\rho_{\phi,*}^{1/4}=\frac{\lambda_{\phi}^{1/4}}{\sqrt{2}}\phi_{*}\ .$ (3.68)
Then, the effective mass of $\Phi$ in the vicinity of the origin is given by
$m_{\phi,\rm eff}^{2}(0)$ $=$ $-\lambda_{\phi}\phi_{0}^{2}+\sum_{i}c_{T}y_{N_{i}}^{2}T_{N,*}^{2}$ (3.69)
$\approx$ $\lambda_{\phi}\phi_{0}^{2}\left[-1+c_{T}\lambda_{\phi}^{5/2}\left(\frac{9}{32\pi}\frac{M_{\rm P}}{\phi_{0}}\right)^{2}\sum_{i}b_{i}\right]$
where $c_{T}=1/12$ [62]. Requiring $m_{\phi,\rm eff}^{2}(0)<0$ to avoid a thermal-trap, we find
$\xi_{\phi}\lesssim 680\left(\frac{0.1}{b_{N}}\right)^{1/5}\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)^{2/5}$ (3.70)
where we took $b_{N}\equiv b_{3}\approx b_{2}\gg b_{1}$. In the case of $\phi_{*}<\phi_{\times}$, the constraint on $\xi_{\phi}$ is much weaker since the would-be temperature of RHNs around the epoch of $\phi=\phi_{\times}$ is much lower. Hence, conservatively speaking, as long as Eq. (3.70) is satisfied, the PQ-field can avoid thermal-trap around the origin, and the PQ-symmetry is never restored during and after inflation.
@@
It should be noted that, even though thermal effect is negligible, and the PQ-symmetry is not restored as a result, domain-wall problem can be avoided only if $\Phi$ has angular motion which is large enough to overcome its quantum fluctuations. Otherwise, $\Phi$ would spread over the whole $2\pi$ angle due to the random scattering of the oscillating field by the barrier around the origin, resulting in the domain-wall problem.
The details of the post-inflation cosmology depends on whether $\phi_{*}$ is larger than $\phi_{\times}$ or not. So, we discuss those cases separately.
Case of $\phi_{*}<\phi_{\times}$:
As shown in Eq. (3.63), $\phi_{*}<\phi_{\times}$ if
$\xi_{\phi}<100\times\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)^{1/2}$ (3.71)
@@
Once $\phi$ decays at the epoch of ${\tilde{H}}_{*}$, the universe will be filled with axi-majorons and RHNs as a comparable or sub-dominant component. Depending on interaction rates, they may or may not reach thermal or kinetic equilibrium among themselves before the standard thermal bath is established from decays of RHNs. Initially, the production of $\phi$ via inverse decay is kinematically forbidden. Hence, thermalization before the decay of RHNs would be driven by interactions of axi-majoron kinetic term, $y_{N_{i}}$, and $y_{\nu_{i}}$. The $\phi$-mediating self-interaction cross-section of axi-majorons is given by
$\sigma_{a,s}=\frac{7s^{3}}{80\pi m_{\phi}^{4}\phi_{0}^{4}}\ ,$ (3.72)
where $s=4E_{a}^{2}$ with $E_{a}$ being the energy of axi-majoron in the center of momentum(CM) frame. If they were not in equilibrium, the number density of axi-majorons would be
$n_{a}(E_{a})=n_{a,*}\left(\frac{E_{a}}{{\tilde{E}}_{a,*}}\right)^{3}$ (3.73)
with
$n_{a,*}\approx\frac{\rho_{\phi,*}}{{\tilde{E}}_{a,*}}$ (3.74)
where $\rho_{\phi,*}\equiv\rho_{\phi}({\tilde{\phi}}={\tilde{\phi}}_{*})$ in Eq. (3.65), and ${\tilde{E}}_{a,*}$ and $E_{a}$ are the energies of an axi-majoron at its production and some time later, respectively. Then, comparing the scattering rate to the expansion rate, one finds that axi-majorons could reach kinetic equilibrium among themselves when its energy is larger than $E_{a,\rm k}$ satisfying
$\frac{E_{a,\rm k}}{{\tilde{E}}_{a,*}}=\left(2^{5}\times\frac{5\pi^{2}}{7}\right)^{1/7}\times\left(\frac{\phi_{0}}{\lambda_{\phi}M_{\rm P}}\right)^{2/7}$ (3.75)
It implies that $E_{a,\rm k}>{\tilde{E}}_{a,*}$ if $\xi_{\phi}\lesssim 100\left(\phi_{0}/\phi_{0}^{\rm ref}\right)^{1/2}$. Hence, for $\xi_{\phi}$ satisfying Eq. (3.71), axi-majorons are never thermalized among themselves after their production in the decay of $\phi$.
@@
where $N=\left(\nu_{R},\nu_{R}^{c}\right)^{T}$ is a four-component Majorana field with the flavor index suppressed, and only relevant leading order terms are shown. When the condition in Eq. (3.56) is satisfied for the RHN-channels of the decay of the inflaton, the cross-sections of the various interaction between the axi-majorons and the RHNs are smaller than the one for axi-majoron self-interactions at high energy limit($s\approx m_{\phi,0}$) at least by a factor of about $b_{i}/3$ (see Appendix C). Moreover, the number density of RHNs is also suppressed by ${\rm B}_{i}$. As a result, for $\xi_{\phi}$ satisfying Eq. (3.71), the associated interaction rates can not catch up the expansion rate for $H<H_{\times}\equiv H(\phi_{\times})$.
Meanwhile, from the ratio of $y_{\nu_{i}}$ to $y_{N_{i}}$, which is given by
@@
one may think that the interaction $y_{\nu_{i}}$ can cause much more efficient scatterings and annihilations of $N_{i}$s to SM particles. However, the decay rate of a RHN mass eigenstate $N_{i}$ is given by (see Appendix D)
@@
where $m_{N_{i}}\gg m_{h}$ was assumed. Also, if the left-handed neutrino mass-eigenstates have the normal hierarchy [15] for their masses as an example, $m_{\nu_{i}}\approx m_{\nu}$ is required for the corresponding heavy RHN mass-eigenstates which we assume to be $N_{2,3}$. Then, for $\xi_{\phi}$ in the range we are interested in, with $\phi_{0}\sim\phi_{0}^{\rm ref}$, heavy RHN mass-eigenstates decay well before their self-interactions/annihilations involving either axi-majorons or SM particles become relevant. The lightest RHNs can also be decoupled from SM particles before their decay if $m_{\nu_{1}}\ll m_{\nu}$ even in the case of $\xi_{\phi}\ll\mathcal{O}(10^{2})$. Therefore, once produced, axi-majorons and RHNs are expected to evolve independently without thermalization at least until the epoch of RHNs decay as long as $m_{\nu_{1}}\ll m_{\nu_{2}}\approx m_{\nu_{3}}\approx m_{\nu}$.
If it were stable, a RHN mass-eigenstate becomes non-relativistic when the expansion rate becomes
${\tilde{H}}_{i,\rm non}={\tilde{H}}_{*}\left(\frac{2m_{N_{i}}}{m_{\phi,0}}\right)^{2}=2b_{i}{\tilde{H}}_{*}\ .$ (3.79)
@@
$\frac{\Gamma_{N_{i}}}{{\tilde{H}}_{i,\rm non}}=\frac{3}{\sqrt{2\lambda_{\phi}}}\frac{m_{\nu_{i}}\phi_{0}}{v_{h}^{2}}\simeq\frac{66}{\xi_{\phi}}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ .$ (3.80)
Also, if stable enough, it would start dominating the universe when the energy of axi-majoron becomes $E_{a}=E_{a,\rm eq}^{i}$ with
$E_{a,\rm eq}^{i}={\rm B}_{i}m_{N_{i}}\ .$ (3.81)